Morning Litho status report 2/4

Gary Sosa gsosa at stanford.edu
Fri Feb 4 07:39:05 PST 2011


Good morning All.....

   Here is the morning Litho status report.

    No airflow in aisle L102(SVG coat/develop aisle) due to transfer fan 
problem. Due to this failure, there will
be a loss of temperature, humidity and particulate control. Repairs are 
in progress. Please consider this
as it relates to the sensitivity of your process. Parts are on order to 
repair the problem

ASML Stepper- System needs 3D Align setup. Also user reports that 
"Dialog Screen" was closed out.
                              Will investigate on morning rounds.

EVG Spray Coater- FSE identified a bad component in the power 
distribution box. Parts are on order
                                    to replace the failed component.

Mask Cleaner- Need to check on status of replacement part.

     Please check back in Coral for latest tool updates and current 
equipment status.

         Thanks.....  The Litho Team


*/Photolithography Area Status/*

	

	

*/Friday, February 04, 2011/*

	

	

	

	

*/General Information / Lab Activities/*

	

	

	

Temp/Humidity/Airflow problem in aisle L-102(SVG Coat/Develop aisle) due 
to transfer fan problem.

Mahnaz- Out today

*/Exposure Tools/*

	

Status

	

Comments

	

Priority

ASML Stepper

	

*PROBLEM*

	

Needs 3D align setup

	

4

EVAlign

	

*UP*

	

	

EVAlign2

	

*UP*

	

	

EVG-Imprint

	

*UP*

	

	

Karl Suss 1

	

*UP*

	

	

Karl Suss 2

	

*UP*

	

	

*/Manual Coaters/*

	

Status

	

	

EVGSpraycoat

	

*DOWN*

	

Problem with power distribution

	

1

Headway

	

*UP*

	

	

Laurel

	

*UP*

	

	

*/Coaters/*

	

Process Status

	

	

SVGCoat

	

3612 Process

	

*Up*

	

	

220-3 Process

	

*Up*

	

	

220-7 Process

	

*Problem*

	

Optimizing Process

	

3

SVGCoat2

	

3612 Process

	

*Up*

	

	

955-0.7 Process

	

*Up*

	

	

3617 Process

	

*Up*

	

	

*/Developers/*

	

Status

	

	

SVGDEV

	

*UP*

	

	

SVGDEV2

	

*UP*

	

	

*/Ovens / Hotplates/*

	

Status

	

	

BLUEM

	

*UP*

	

	

Blue M 90 C.

	

*UP*

	

	

Blue M 110 C

	

*UP*

	

	

Blue M 150 C

	

*UP*

	

	

Fusion

	

*DOWN*

	

Hotplate experiments

	

UV-Cure

	

*UP*

	

	

White-Oven

	

*UP*

	

	

YES Vapor Prime

	

*UP*

	

	

*/Hotplates- /*

	

	

	

Laurel-left

	

*UP*

	

	

Laurel- right

	

*UP*

	

	

Headway

	

*UP*

	

	

Torrey Pines- left

	

*UP*

	

	

90 Deg C

	

*UP*

	

	

115 Deg C

	

*UP*

	

	

Torrey Pines- Right

	

*UP*

	

	

*/Metrology/Inspection/*

	

Status

	

	

Alpha-Step 500

	

*UP*

	

	

Nanospec 2-Litho

	

*UP*

	

	

Nanospec 1-Etch

	

*UP*

	

	

P2 Profiler

	

*PROBLEM*

	

-11 stage error

	

7

*/Microscpoes-/*

	

	

	

#1 Leica/Analytical

	

*UP*

	

	

#2 Ergolux/Drytek2

	

*UP*

	

	

#3 Olympus/Litho

	

*UP*

	

	

#4 Nikon Optiphot/Litho

	

*UP*

	

	

#5 Nikon/Litho

	

*UP*

	

	

#6 Leitz/Bosch

	

*UP*

	

	

#7 Leitz/Diffusion

	

*UP*

	

	

#8 Olympus/Litho

	

*UP*

	

	

#9 Olympus/Litho

	

*UP*

	

	

*/Miscellaneous/*

	

Status

	

	

EVBond

	

*PROBLEM*

	

Left flag did not pull

	

5

KSBonder

	

*PROBLEM*

	

Leaking nitrogen hose

	

6

KSCoat

	

*UP*

	

	

Litho SRD

	

*UP*

	

	

Mask Cleaner

	

*DOWN*

	

Water leak, Waiting for parts

	

2

	

	

	




-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/litho/attachments/20110204/f3b359eb/attachment.html>


More information about the litho mailing list